# LLaVA-MORE [Uncaptioned image]: A Comparative Study of LLMs and Visual Backbones for Enhanced Visual Instruction Tuning

### v1

@@ 4.4 Changing the Visual Backbone
The only exception is MME-C, where DINOv2 with registers scores the highest at the 3.8B scale, and DINOv2 is the best at the 9B scale with 334.3 points. Among the models who enjoyed image-text pre-training, SigLIP greatly improves over CLIP at both scales, even though it forces the LLM to process 26% more visual tokens. Despite its more complicated training recipe, the new SigLIP2 performs on par with the original SigLIP at the 3.8B scale but records an average 0.4% gain over SigLIP at the 9B scale. For this reason, we will always include SigLIP2 as the visual backbone in all the subsequent experiments.
[…]
4.5 Changing the Image Resolution
The choice of a robust visual backbone and the appropriate LLM are undeniably critical factors that significantly influence the overall performance of a multimodal model. However, an equally important factor, is the resolution of the input image, which plays important role in visual understanding. Higher image resolutions provide more fine-grained visual information, which can significantly aid the MLLM in better interpreting the content and thereby improving the overall performance.
@@ 4.5 Changing the Image Resolution
Notably, as shown in Table 3 (top), it is evident that S2 generally enhances the performance of LLaVA-MORE-3.8B when CLIP is used as the visual backbone. The improvements become even more consistent upon switching to SigLIP2, which operates at a higher resolution than CLIP. However, when scaling up to LLaVA-MORE-9B (Table 3 bottom), the benefits of S2 vanish for certain benchmarks. For instance, on Science-QA and AI2D, S2 appears detrimental at the 9B scale, while it is advantageous with LLaVA-MORE-3.8B.
From these results, we can conclude that small-scale MLLMs may greatly benefit from working with high-resolution images. However, as the model scales up, the positive impact of increased resolution starts to diminish, indicating that the benefits of higher resolution are not universally applicable and depend on the specific task at hand.
@@ 4.6 Dataset Contribution During Pre-Training
Interestingly, the results highlight that LLaVA-MORE-3.8B never records the highest score when using the original mixture. Conversely, exclusively training on LAION stands out as the winning choice in 8 out of 13 benchmarks. A possible explanation for that is that LAION consists of noisy image-text pairs collected from the web, that closely resemble the nature of the training data used by the SigLIP2 visual encoder that powers LLaVA-MORE-3.8B. LLaVA-MORE-9B demonstrates greater robustness to variations in the pre-training dataset, with the highest score being more spread across the different configurations. Curiously, Recap samples enhance the Chinese fluency of LLaVA-MORE-9B, leading to the best score of 69.4 on MMB-Cn, and generally improve the results on other MLLM benchmarks.
@@ 5 Conclusion
Acknowledgments. We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources. This work has been conducted under two research grants, one co-funded by Leonardo S.p.A. and the other co-funded by Altilia s.r.l., and supported by the PNRR-M4C2 project “FAIR - Future Artificial Intelligence Research” and by the PNRR project “ITSERR - Italian Strengthening of Esfri RI Resilience” (CUP B53C22001770006), both funded by the European Union - NextGenerationEU.

### v2

@@ 4.5 Changing the Image Resolution
Figure 3: Qualitative of image descriptions generated by three MLLMs (LLaVA-1.5-7B, LLaVA-MORE-3.8B, and LLaVA-MORE-9B). The example shows differences in detail, context, and narrative style across scenarios like sports, transport, toys, and portraits.
@@
Notably, from Table 3 (top), it can be seen that S2 generally improves LLaVA-MORE-3.8B when using CLIP as the visual backbone. The improvements are even more consistent when switching to SigLIP2, which already works at a higher resolution than CLIP. However, when scaling up to LLaVA-MORE-9B (Table 3 bottom), the benefits of S2 vanish for some benchmarks. For instance, on Science-QA and AI2D, S2 appears detrimental at the 9B scale, while it is advantageous with LLaVA-MORE-3.8B.
From these results, we can conclude that small-scale MLLMs may greatly benefit from working with high-resolution images. However, the positive impact of higher resolution appears to diminish as model size increases, and, ultimately, the benefits of increasing image resolution seem to be highly task-dependent. For instance, with TextVQA, a benchmark that heavily relies on detecting and recognizing text within images, S2 consistently brings improvements. A similar pattern is found on GQA, which challenges MLLMs with questions requiring in-depth scene understanding. Curiously, we find substantial gains by applying S2 on the Chinese version of MMBench (MMB-Cn), especially at the 3.8B scale, while its behavior on the English version is conflicting: S2 improves with SigLIP2, but it is detrimental with CLIP.
@@ 4.6 Dataset Contribution During Pre-Training
Interestingly, LLaVA-MORE-3.8B never achieves the top score with the original mixture. Conversely, exclusively training on LAION stands out as the winning choice in 8 out of 13 benchmarks. This may be because LAION’s noisy web-sourced pairs closely resemble the nature of the training data used by the SigLIP2 visual encoder behind LLaVA-MORE-3.8B. In contrast, LLaVA-MORE-9B shows more robust performance across the different configurations. Notably, adding Recap samples boosts LLaVA-MORE-9B’s Chinese fluency, leading to the best score of 69.4 on MMB-Cn and improving other MLLM results.
@@
4.7 Qualitative Results
To better understand the behavior of different MLLMs beyond their quantitative performance on standard MLLM and VQA benchmarks, we also present qualitative results on the image captioning task [^1] For these qualitative results, we use the prompt: “Describe this image.”., as shown in Fig. 3. In particular, we compare LLaVA-1.5-7B [42], which serves as our baseline, against our model based on Phi-4-3.8B and Gemma-2-9B using SigLip2 as visual encoder. As it can be seen, LLaVA-MORE versions can effectively describe input images and provide detailed and rich textual descriptions. For example, only the two versions of LLaVA-MORE successfully avoid hallucinations and provide an accurate description of the scene. In contrast, the LLaVA-1.5-7B model incorrectly identifies the object in question (i.e., erroneously recognizing the snowboard in the image as skis) highlighting a notable failure in visual understanding and alignment with textual output.
@@ 5 Conclusion
Limitations
In this work, an analysis of a subset of publicly available LLMs and visual backbones was considered. The rapid evolution of this research area means that the results presented provide a snapshot of the best configurations currently achievable. In addition, preliminary results are presented on the integration of text-only reasoning models into multimodal tasks through fine-tuning. These insights can serve as a starting point to bridge the gap between text-only and multimodal reasoning.
Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources. This work has been conducted under two research grants, one co-funded by Leonardo S.p.A. and the other co-funded by Altilia s.r.l., and supported by the PNRR-M4C2 project “FAIR - Future Artificial Intelligence Research” and by the PNRR project “ITSERR - Italian Strengthening of Esfri RI Resilience” (CUP B53C22001770006), both funded by the European Union - NextGenerationEU.